### v1

# Models of attractor dynamics in the brain

Abstract
Attractor dynamics are a fundamental computational motif in neural circuits, supporting diverse cognitive functions through stable, self-sustaining patterns of neural activity. In these lecture notes, we review four key examples that demonstrate how autoassociative neural network models can elucidate the computational mechanisms underlying attractor-based information processing in biological neural systems performing cognitive functions. Drawing on empirical evidence, we explore hippocampal spatial representations, visual classification in the inferotemporal cortex, perceptual adaptation and priming, and working-memory biases shaped by sensory history. Across these domains, attractor network models reveal common computational principles and provide analytical insights into how experience shapes neural activity and behavior. Our synthesis underscores the value of attractor models as powerful tools for probing the neural basis of cognition and behavior.
Keywords attractor dynamics autoassociative neural networks neural dynamics
Introduction
The brain’s remarkable computational abilities emerge from the complex interplay of neural circuits organized into distinct, yet interconnected, functional architectures. Among these architectural motifs, canonical microcircuits with recurrent connectivity patterns represent fundamental computational units that appear across diverse brain regions (Douglas et al., 1989; Bastos et al., 2012). These recurrent connections, where neurons form closed loops of excitation and inhibition, are ubiquitous throughout cortical and subcortical structures and play a crucial role in information processing beyond what purely feed-forward architectures could achieve (van Bergen and Kriegeskorte, 2020).
Recurrent neural networks support a class of dynamics known as attractor dynamics, where network activity evolves toward and stabilizes around specific patterns—attractors in the state space of possible neural activations (Hopfield, 1982; Amit, 1989). These attractors can manifest as point attractors (stable equilibrium states), line or ring attractors (continuous manifolds of stable states), or limit cycles (periodic trajectories), providing the substrate for persistent neural activity essential for working memory, perception, decision-making, spatial navigation, and planning. The computational richness of attractor networks stems from their ability to store multiple stable states, implement pattern completion from partial inputs, resist noise, and support information integration over time (Wang, 2001; Rolls, 2007; Tang et al., 2018)
The expressive power of attractor dynamics as models of biological neural computation has received substantial empirical support across a range of cognitive domains. In this review, we summarize two lectures delivered by Dr. Athena Akrami at the 2023 School on Analytical Connectionism—an event focused on analytical tools for probing neural networks and higher-level cognition—held at University College London. The lectures illustrate how attractor-based models can illuminate core neural computations across systems. We begin with bifurcation phenomena in hippocampal place cell remapping, followed by attractor dynamics in inferotemporal cortex during visual classification. We then examine the interplay of adaptation and priming effects in perception, and conclude with coupled attractor dynamics between cortical regions that give rise to working memory biases. Throughout, we highlight how recurrent neural computations and their associated attractor dynamics offer a unifying computational framework for understanding seemingly disparate neural and cognitive phenomena.
1 Attractor dynamics in hippocampal place cells
The hippocampus is widely regarded as a hotspot for investigating attractor dynamics, largely due to its central role in encoding spatial representations and memory retrieval (Knierim and Zhang, 2012; Whittington et al., 2020). In particular, recurrent connectivity within the CA3 region of the hippocampus is believed to implement auto-associative network dynamics, capable of maintaining stable attractor states that support robust spatial representations observed in CA1 (Rolls, 2007). Hippocampal place cells, which exhibit highly selective firing patterns corresponding to specific locations in space, offer an ideal test-bed for evaluating principles of attractor-based computation and theoretical models grounded in these dynamics.
Wills et al. (2005) investigated hippocampal attractor dynamics by examining changes in place cell firing patterns as rats explored environments of varying shapes. Rats were first familiarized over six days with two distinct enclosures, one square and one circular, during which hippocampal place cells in CA1 exhibited global remapping between the two shapes in four out of six animals (Figure 1A). On the seventh day, environments were systematically morphed from square to circle in intermediate (octagonal) shapes (square-like and circle-like morph trials were alternated), and the response of individual place cells was monitored. If the place cell firing patterns exhibit attractor dynamics, then the patterns for the square-like shapes should closely resemble the firing pattern in the square enclosure, and the circle-like shapes should closely resemble the firing pattern in the circular enclosure. Indeed, cells demonstrated an abrupt, attractor-like switch between square-like and circle-like firing patterns (Figure 1B,C), observable even within the initial seconds of each morph trial, although these transitions became increasingly pronounced with continued exposure. These findings provide strong empirical support for attractor dynamics in hippocampal spatial representations. Interestingly, contrasting results from Leutgeb et al. (2005) showed more graded transitions in similar morphing paradigms, suggesting that the hippocampus can operate in different dynamical regimes—discrete or continuous—depending on factors such as input ambiguity, learning history, and task demands (Tsodyks, 2005).
Subsequent studies further refined this view. Colgin et al. (2010) demonstrated that abrupt global remapping in the hippocampus cannot be fully accounted for by direct, feature-based differences between environments, as simpler autoassociative models would suggest. Instead, this remapping reflects the activation of distinct path-integration reference frames shaped by environmental features, such as geometry. These reference frames are generated primarily by the medial entorhinal cortex (Hafting et al., 2005), which provides spatial input to the hippocampus. Together, these results underscore how attractor dynamics within the hippocampal–entorhinal circuit shape spatial memory representations, with environmental geometry and path integration playing critical roles in anchoring these neural cognitive maps.
2 Attractor dynamics in inferior temporal cortex
The inferior temporal (IT) cortex is critical for visual object recognition (DiCarlo et al., 2012) and long-term visual memory storage (Sakai and Miyashita, 1991). Given its rich recurrent connectivity and involvement in memory retrieval, IT has been hypothesized to exhibit attractor dynamics similar to those observed in the hippocampus. Attractor network models have been proposed as computational mechanisms to support categorization and stabilize object representations, particularly under conditions of visual ambiguity (Miyashita et al., 1993; Rolls, 2007).
In their study, Akrami et al. (2009) directly tested this hypothesis, probing attractor-like categorization dynamics in IT using a visual morphing paradigm conceptually analogous to the hippocampal place cell studies discussed earlier. Monkeys performed a match-to-sample task, discriminating between pairs of familiar photographic stimuli (“endpoints”) and intermediate morphs generated via nonlinear pixel-wise blending (Figure 2A). The match options were always a pair of endpoints, corresponding to the sample, which could be one of the endpoints or their morph. Single-electrode recordings in anterior IT cortex (area TE and adjacent perirhinal cortex) targeted neurons selectively responsive to one of the endpoint images (designated “effective”) but not its paired counterpart (“ineffective”). Early neural responses (-ms after stimulus onset) scaled linearly with stimulus similarity to endpoints, encoding the morph level (Figure 2B). However, during a later response period (–ms post-stimulus), firing rates for morph stimuli similar to the effective stimulus showed convergence, losing their linear dependence and approaching the firing rate elicited by the effective endpoint. This convergence was asymmetric: morphs closer to the ineffective endpoint maintained linearly graded responses, suggesting a selective attractor basin biased toward the effective memory. Notably, the strength of this asymmetric convergence grew with the animals’ behavioral proficiency, indicating experience-dependent shaping of attractor-like dynamics in IT.
To explore the underlying mechanisms, Akrami et al. (2009) implemented an autoassociative neural network model (Figure 2C) comprising two layers of neurons each: an input layer mimicking early visual representations and a recurrent output layer simulating the IT cortex. Each output neuron received sparse, randomly assigned feedforward inputs from neurons in the input layer and recurrent inputs from randomly selected neurons within the output layer.
Visual memory patterns were stored within the recurrent connections using a Hebbian covariance learning rule:
where is the synaptic weight between neurons and , indicates the presence (1) or absence (0) of a connection from neuron to neuron , is the activity of neuron in pattern , is the mean activity across all patterns, is the number of recurrent connections per neuron, and is the activity sparseness parameter.
To simulate memory storage of familiar visual stimuli, a set of random but structured activity patterns (“stored patterns”) was first established. These consisted of sparse firing rate vectors across the input-layer neurons, drawn independently from a truncated logarithmic distribution. Each stored pattern corresponded conceptually to one familiar (“endpoint”) visual stimulus used in the experiment. Morph stimuli were then created by systematically blending pairs of stored patterns—analogous to intermediate morphs used in the empirical study—by replacing the firing rate of a randomly selected subset of neurons in one stored pattern with the corresponding values from another. Importantly, the network stored only the endpoint patterns, not the intermediate morphs, consistent with the assumption that monkeys would not form stable memory traces for ambiguous intermediate stimuli that lacked distinct behavioral relevance.
The model included spike-frequency adaptation to mimic firing rate decay observed in cortical neurons. Adaptation was implemented by subtracting a term proportional to each neuron’s recent activity from its total synaptic input, as:
with:
Here, is the activity of neuron at time , is the summed synaptic input, is the firing threshold, is the neural gain, and parameters , , and control the strength and time-scale of adaptation.
Simulations revealed that asymmetric attractor-like convergence of neural responses emerged specifically when morph stimuli were interpolated between a stored (“effective") and an unstored (“ineffective”) pattern, and the network operated near its maximum memory storage capacity (approximately patterns in this model; Figure 2D). Under these conditions, responses to morphs resembling the stored pattern converged toward the corresponding attractor state, while responses to morphs closer to the unstored pattern remained linearly graded (i.e., dependent on morph level). In contrast, when memory load was low (e.g., patterns), convergence was overly broad with almost all morphs attracted to the stored memory, inconsistent with sharp behavioral categorization and neural responses observed empirically. Furthermore, when both morph endpoints were stored, convergence occurred symmetrically, eliminating the observed in IT responses. To explain this, the authors proposed that this discrepancy could reflect the experimental selection bias in neuron sampling: neurons were selected based on their responsiveness to one of the stimuli, likely belonging to the attractor basin of the "effective" pattern, whereas the "ineffective" pattern may have been represented elsewhere in IT (Haxby et al., 2001; Kiani et al., 2007). Thus, the observed asymmetry likely reflects this experimental bias in selective sampling of neuronal populations belonging predominantly to one endpoint’s memory representation rather than a fundamental computational difference between stored and unstored patterns. Finally, the inclusion of spike-frequency adaptation improved the model’s match to experimental data by allowing network activity to decay over time after stimulus offset and effectively replicating the experimentally observed temporal response dynamics.
Overall, the modeling results support the view that category-specific convergence in IT emerges from attractor dynamics within local recurrent networks, shaped by memory load and stimulus familiarity. This framework provides a mechanistic account of how experience-dependent plasticity gives rise to categorical visual representations in the IT cortex.
3 Prior experience and perceptual biases
Perception is not a passive reflection of sensory input, but a constructive process profoundly shaped by prior experience that gives rise to systematic perceptual biases (Helmholtz, 1924; Beck, 1967). When faced with ambiguous stimuli, the brain actively resolves uncertainty by integrating contextual information and drawing on memory-based predictions (Friston, 2005; Bar, 2007). Prior exposure to prototypical stimuli can bias perception in two seemingly opposing directions: adaptation aftereffects, where perception is repelled away from the recently experienced adapter stimulus, and priming effects, where perception is attracted toward it (Logothetis and Pauls, 1995; Tulving and Schacter, 1990). A compelling example occurs in facial emotion perception, where prolonged exposure to a happy face causes a subsequent neutral face to appear slightly angry—a classic repulsive aftereffect (Webster et al., 2004; Aguado et al., 2007), as shown in Figure 3. However, a brief exposure to a happy face can cause subsequent faces to be perceived happier-a classic priming effect (Murphy and Zajonc, 1993).
Several theoretical frameworks have attempted to reconcile these effects by appealing to differences in temporal dynamics, neural substrates, or functional roles. Early formulations, such as adaptation-level theory in Helson (1964), modeled how prior experiences set perceptual reference points. Later, following Barlow’s theory of sensory recalibration and predictive coding frameworks (Barlow, 1993), these phenomena have been recast within a Bayesian inferential model, in which perception results from the integration of sensory evidence and prior expectations (Kersten et al., 2004; Fritsche et al., 2017).
In a different take, Akrami et al. (2010) offered a unified mechanistic explanation using attractor neural network models incorporating firing rate adaptation. In this framework, the same visual input can produce either adaptation or priming depending on the temporal dynamics of network activity and the stability of stored memory representations. Specifically, short-lived inputs may nudge activity toward a familiar attractor, yielding priming, while sustained stimulation can destabilize that attractor via adaptation, producing repulsion. This model provides an account of how opposing perceptual biases emerge from a single underlying circuit architecture modulated by memory and experience (see below for details).
From Adaptation to Priming: The Role of Task Delays
Adaptation aftereffects represent well-documented phenomena in vision research (Webster and Maclin, 1999; Clifford et al., 2007), and while traditionally studied with low-level features such as color and orientation (Gibson, 1937; Blakemore and Campbell, 1969), these effects extend to high-level domains including face (Leopold et al., 2001; Rhodes et al., 2003) and complex object (Daelli et al., 2010) perception. However, experiments that explore both adaptation and priming effects within a single paradigm, particularly for non-face objects, have been poorly documented.
Building on the attractor-based models of perception and categorization of morphed stimuli discussed in Section 2, Daelli et al. (2010) investigated how perceptual adaptation influences the interpretation of ambiguous real-world object images and how these perceptual biases evolve. Crucially, their study systematically explored how task parameters-such as the duration of the adapting stimulus, the characteristics of the test image, and the delay interval between the two-modulate perceptual outcomes.
The authors conducted three behavioral experiments using morphed images of animals, plants, and objects under varying temporal conditions (Figure 4A). In the first experiment, participants were presented with a clear prototype image (adapter) followed shortly (ms delay) by a morph between that prototype and another object. Participants consistently exhibited a repulsive bias, perceiving the morph more dissimilar to the adapter than it actually was (Figure 5-top), demonstrating adaptation aftereffects in the perception of complex, non-face objects. To establish the temporal stability of these adaptation aftereffects, the second experiment introduced a longer delay between adapter and target (ms). Surprisingly, the repulsive effect disappeared and was replaced by an attractive priming effect—participants were now more likely to judge the ambiguous morph as resembling the adapter (Figure 5-bottom). This temporal reversal suggests that adaptation effects weaken over time, revealing a slower-developing priming mechanism that biases perception toward previously seen stimuli. In the final experiment, the researchers tested whether adaptation could still occur when the adapter itself was ambiguous. Unlike in previous experiments, adaptation to an ambiguous stimulus consistently led to priming, regardless of the duration of the delay. This finding suggests that adaptation requires a well-defined, strongly encoded prototype to exert a repulsive influence on perception, while ambiguous stimuli are more likely to produce attractive biases.
Collectively, these findings demonstrate that the same stimuli and task can elicit either adaptation or priming effects, depending on temporal task parameters, highlighting the dynamic nature of perceptual biases. The results point to two opposing yet interrelated neural mechanisms: adaptation, which induces repulsive aftereffects by attenuating responses to recently encountered stimuli, and memory-based priming, which attracts perception toward prior stimuli via the activation of stable memory representations. When a stimulus is well-defined and recently presented, adaptation dominates, pushing perception away. As time passes, or when adaptation is weak (as with ambiguous stimuli), priming emerges, pulling perception toward familiar experiences. This work offers a compelling demonstration of how perceptual history modulates ongoing experience and provides key insights into the temporal dynamics that govern the balance between adaptation and memory-based priming.
A unifying attractor dynamics model to short-term visual experience
The studies reviewed above showed that perception is inherently contextual, relying on the integration of prior experience with incoming sensory input to dynamically resolve categorical decisions. This integration is fundamentally rooted in the temporal dynamics of neurons involved in perceptual computations. To address this, Akrami et al. (2010) developed a neural network model that unifies historically separate perceptual phenomena-adaptation aftereffects and priming-within a single theoretical framework grounded in attractor dynamics. Their model reproduces key electrophysiological and behavioral findings from Daelli et al. (2010) and accounts for a range of priming effects observed in humans.
The network architecture is an auto-associative memory model comprising recurrently connected neurons with threshold-linear activation functions. Crucially, it incorporates firing rate adaptation, a biophysical property of pyramidal neurons, which allows the network to transition between transient and stable states. The investigation centers on the interplay between recent input-driven activity and stable attractor dynamics to show how this interplay gives rise to systematic perceptual biases.
To quantify these effects, the model defines an index measuring the distance between two network states at any time:
where is the overlap with pattern A when A is the adapter, and is the overlap with A when B is the adapter. This metric allowed the authors to systematically map the influence of adaptation strength, delay, and target duration on perceptual outcomes across a broad range of parameter space.
The simulations revealed that firing rate adaptation is the essential component for producing adaptation aftereffects (Figure 6B), while two distinct mechanisms underlying priming effects were identified:
Perceptual Bias (Type A Priming): When the adapter induces an attractor state and the target provides a weak or ambiguous input, the network remains in the prior state, biasing perception toward the adapter. This occurs primarily when adaptation is minimal (Figure 6A).
Double Adaptation (Type B Priming): Under stronger firing rate adaptation, the network can undergo multiple transitions between attractor states. Under specific temporal conditions, this results in the network settling back into the attractor associated with the adapter, paradoxically producing a priming effect even in the presence of adaptation. (Figure 6C).
The research also examined backward masking effects, where a secondary stimulus disrupts the processing of the initial adapter. Simulations revealed that if a mask is presented before the network fully settles into an attractor state, adaptation aftereffects are substantially weakened, leaving predominantly priming effects. This provides a mechanistic explanation for previously observed interactions between masking and adaptation.
By systematically varying parameters—including memory load (i.e., number of stored patterns), firing rate adaptation strength, and the nature of adapter stimulus (ranging from well-defined prototypes to ambiguous morphs)—the model maps how network dynamics shift between adaptation aftereffects and priming regimes. Robust adaptation effects were shown to require strong firing rate adaptation, while priming emerged in networks with stable attractor representations, contingent on the timing of the adapter and target presentation.
The temporal relationship between stimuli proved critical. Brief adapter-target intervals favored adaptation aftereffects, while extended delays promoted priming. Moreover, target duration emerged as an additional factor modulating outcomes: longer target exposures could reverse expected effects by allowing the network to overcome adaptation and converge on a stored attractor. The model’s predictions align closely with behavioral findings from face and motion adaptation studies, reinforcing the model’s relevance to real-world perceptual dynamics.
These findings underscore the importance of firing rate adaptation as a key modulator of perceptual dynamics and highlight how attractor networks provide a mechanistic basis for the flexible, experience-dependent biases observed in human perception.
4 Coupled PPC and PFC Dynamics Underlying Biases in Working Memory
Behavior in perceptual and working memory tasks is often shaped by past sensory experiences. Two prominent forms of such biases are the contraction bias- where working memories of stimuli are pulled toward the long-term average of past inputs (Jazayeri and Shadlen, 2010; Raviv et al., 2012)-and the recency bias, or serial dependence, where recent stimuli exert a disproportionate influence on current judgments (Fischer and Whitney, 2014; Barbosa et al., 2020). Although traditionally viewed as distinct phenomena arising from different cognitive mechanisms, recent evidence suggests they may share a common neural substrate. Notably, Akrami et al. (2018) showed that silencing the posterior parietal cortex (PPC), in rats, significantly reduces both contraction and recency biases, implicating this region in integrating sensory history. Building on these findings, Boboeva et al. (2024) proposed a mechanistic model showing how interaction between the PPC and a downstream working memory (WM) area (e.g., prefrontal cortex, though not proven yet) could give rise to these working memory biases. In what follows, we examine these two studies in detail, focusing on how PPC-WM dynamics may underlie sensory-history-dependent distortions in working memory.
To explore these biases experimentally, Akrami et al. (2018) utilized a parametric working memory (PWM) task, involving sequential presentation of two graded stimuli separated by a delay interval. In their auditory version of the task (illustrated in Figure 7A), rats are presented sequentially with two tones, and , and must determine which tone was louder, thereby engaging working memory processes rather than immediate sensory comparisons. This delay interval provides a window for prior sensory experience to bias the internal representation of .
A key aspect of the PWM task is the manipulation of stimulus pairs (presented at each trial) based on their proximity to the identity line = (Figure Figure 7B), which changes the comparison difficulty. Pairs closer to the diagonal line are harder to discriminate due to smaller intensity differences. Rats were presented with stimulus pairs that were all equally distant from the diagonal (see Figure 7B). Despite the equal objective difficulty, performance was systematically modulated by contraction bias: the remembered value of appeared to be shifted toward the average of previously encountered stimuli (see Figure 7A). When this shift increased the perceived distance from (bias+), performance improved; when it reduced the difference (bias-), performance declined.
To quantify these history affects, Akrami et al. (2018) used a logistic regression to model behavioral choices as a function of current stimulus values , the average of past stimuli (capturing contraction bias), the stimulus pair from the immediately preceding trial (capturing recency), and reward and choice history. This analysis revealed clear evidence for both contraction and recency biases. Importantly, similar biases were observed in human participants, suggesting conserved underlying mechanisms.
To establish a causal role for PPC, optogenetic inactivation experiments were conducted. Temporary silencing of PPC during the delay period abolished both biases (Figure 7C), without impairing overall task performance, indicating that PPC specifically contributes to integrating sensory history into working memory representations rather than maintaining WM information per se. Additionally, the PPC was shown to encode sensory history (Figure 7D).
Building on this, Boboeva et al. (2024) developed a computational model comprising two interacting line-attractor networks (Figure 7E): a slowly integrating PPC network with adaptive dynamics and a WM network downstream to PPC with more stable, persistent activity that responds to sensory inputs with fast dynamics. Both networks receive direct sensory input, with the PPC accumulating sensory information over time and projecting to the WM, thereby modulating its memory representations.
Sensory-history biases naturally emerge from differences in integration timescales between the PPC and WM networks. During the delay interval, the memory representation in the WM network either remains stable or is shifted by PPC input, introducing bias into working memory. See Figure 7F for a sample trial. Crucially, the model predicts stronger biases with increased delay intervals, aligning with experimental observations. Moreover, when PPC input was removed from the model, the biases disappeared, mirroring the inactivation findings from Akrami et al. (2018) and reinforcing the PPC’s central role as a sensory history integrator.
To complement the mechanistic account, Boboeva et al. (2024) also provided a probabilistic model, summarizing the network dynamics. In this simpler framework, the memory of either remains veridical (unchanged with high probability) or is resampled from a distribution reflecting past stimuli. Despite its simplicity, this model captured key behavioral patterns and mirrored predictions of the neural network model, providing convergent evidence that history-dependent biases in working memory can emerge from probabilistic inference shaped by PPC-WM neural dynamics.
Together, these studies underscore a critical role for PPC-WM interactions in generating working memory biases driven by sensory history. They provide a compelling framework linking behavioral phenomena such as contraction and recency biases to specific neural mechanisms, integrating experimental, computational, and theoretical perspectives on memory distortions.
Conclusion
Taken together, the four examples reviewed in these lectures highlight the utility of auto-associative neural networks as a powerful and versatile class of analytical models for studying attractor dynamics—a computational motif that recurs across diverse brain regions and cognitive domains. While these models are intentionally simplified and may not capture the full complexity of neural responses to naturalistic stimuli or real-world behavioral tasks, their strength lies in their analytical tractability. By enabling researchers to precisely isolate, manipulate, and interpret specific features of neural dynamics, these models yield mechanistic insights that are often obscured in more complex, high-dimensional systems. Importantly, auto-associative networks provide a conceptually transparent bridge between neural activity and behavior, revealing how stable patterns of activity can support memory, perception, categorization, and decision making. These insights complement the growing class of large-scale, image-computable neuroconnectionist models that prioritize biological realism and scale Doerig et al. (2023).
While future work must address how to extend attractor-based frameworks to operate in more realistic, high-dimensional sensory spaces (Goetschalckx et al., 2023; Thorat et al., 2023; Soo et al., 2024), the foundational principles uncovered by analytical models remain essential. Their simplicity is not a limitation but a strength, offering a conceptual framework for hypothesis testing, theory building, and ultimately, for understanding the computational architecture of the brain.